EG012

PI: Gurvits
Title: High Z quasars

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: B
Obs time (UT): 299 EU 17:30(25/10)-10:00(26/10)
Stations: EbJbNtToShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Torun (Roman Feiler at Sat Oct 26 10:13:31 EDT 1996) Success
Part 1 - no known problems

Effelsberg (D.Graham at Tue Oct 29 02:20:13 EST 1996) Failure
Approximately halfway through the experiment, the operator tried to improve focus. He entered a value with wrong sign, so that focus was then wrong by 10cm, reducing antenna gain almost to zero. 50% of experiment lost.

HartRAO (Jonathan Quick at Tue Oct 29 04:56:56 EST 1996) Minor failures
LO mis-set from 21:55 thru 22:57 - no other problems to report

Westerbork (tony foley at Tue Oct 29 09:06:49 EST 1996) Success
looks OK

Noto (Corrado Trigilio at Wed Oct 30 05:51:58 EST 1996) Success
No problems reported. Cloudy sky.

Shanghai (Huang Xinyong at Tue Nov 5 03:46:53 EST 1996) Minor failures
The data is invalid from 17:32:30-18:01:04, because the tape vacuum failed. Weather was overcast.

Jodrell (Alastair Gunn at Thu Nov 7 07:00:46 EST 1996) Minor failures
No known problems. Weather good. Tapes shipped to Bonn on 28/10/96 via Woodcroft.

Urumqi (Liu Xiang at Mon Nov 18 08:53:47 EST 1996) Success
Clock offset 0.0+/-0.5us.

Simeiz (Alexandre Volvatch at Tue Nov 19 03:53:07 EST 1996) Success
No problems.

MPIfR (Scott Aaron at Tue Sep 23 14:28:20 METDST 1997) Minor failures
Standard Westerbork problem. Rest OK.